OpenStudy (anonymous):

f^-1(3) when f(x) = 2x + 3 / 5

OpenStudy (anonymous):

\[f ^{-1}(3) when f(x) \frac{ 2x+3 }{ 5 }\]

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

you want to know for what value of x, the function spits out 3

ganeshie8 (ganeshie8):

so set the function equal to 3 and solve x

ganeshie8 (ganeshie8):

\[\large 3 = \dfrac{2x+3}{5}\] solve \(x\)

OpenStudy (anonymous):

so x = 6

ganeshie8 (ganeshie8):

Yep!

OpenStudy (anonymous):

is this right? f(1) when f(x) = 5x + 1 f(x) = 5(1) + 1 x = 6

ganeshie8 (ganeshie8):

right!

OpenStudy (anonymous):

and last part of the question 3y - 7 = y + 5 -5 - 5 3y - 12 = y y = 6

ganeshie8 (ganeshie8):

Excellent! so all 3 parts have the same answer of 6 is it

OpenStudy (anonymous):

Yea it was asking Which of the following statements have the same result? Explain each step in solving each one.

ganeshie8 (ganeshie8):

Makes sense!

OpenStudy (anonymous):

ok thanks once again! :)

ganeshie8 (ganeshie8):

np :)